# Not All Attention Heads Are What You Need: Refining CLIP’s Image Representation with Attention Ablation

### v1

@@ 5.1 Experimental Settings & Implementation Details
Table 4: Comparison on Flickr30k for retrieval.
@@
tasks ACT CEL COL CNT FIG OBJ OCR POS SCE mean CR NC TT mean
@@
Table 5: Text-to-image retrieval on ReCoS based on ViT-B, evaluated by R@1.
5.2 Main Results
5.2.1 AAT-improved CLIP for Retrieval
We evaluate CLIP models improved using both AAT-GA and AAT-BP, reporting retrieval performance across different sizes, including ViT-B, ViT-L, and ViT-H. Results are presented for MS COCO, Flickr30k, and ReCoS, with comparisons against the baseline vanilla CLIP.
@@ 5.2.1 AAT-improved CLIP for Retrieval
Table 3 presents the results on MS COCO, achieved by AAT-GA and AAT-BP for text-to-image and image-to-text retrieval. For text-to-image retrieval, GA and BP perform comparably, with mean-R differences under 1%. On the test set, AAT yields a 1.3$\sim$2.6gains in mean-R across model sizes, and over 1.5% on average for the all set. For image-to-text retrieval, AAT achieves a 0.8$\sim$2.1mean-R gain on the test set and 0.6$\sim$2.7on the all set.
@@
Table 7 reports results on Flickr30k, showing a 1.08$\sim$1.93mean-R improvement on the test set and 1.64$\sim$2.56on the all set for text-to-image retrieval using AAT-improved CLIP. Notably, gains on the test set tend to diminish at higher recall levels, likely due to saturation. A similar trend is observed for image-to-text retrieval as well, yet with consistent mean-R improvements of 0$\sim$0.8on the test set and 1.08$\sim$3.61on the all set.
ReCoS
To demonstrate the broader effectiveness of AAT, Table 5 provides text-to-image retrieval results on the challenging ReCoS, spanning 12 diverse subsets across various domains (see Appendix A for subset abbreviations). These subsets are grouped into two categories based on image domains: natural and non-natural. The natural includes real-world imagery, such as landscapes, animals, movie scenes, and profile photos. In contrast, the non-natural comprises synthetic visuals like code snippets, characters, and formulas, typically displayed on white or gray backgrounds.
AAT consistently outperforms on natural-scene subsets, with a mean R@1 gain of around 3%. However, its performance declines on non-natural images. We attribute this drop to two main factors: (1) significant domain shifts between the test images and those used to identify detrimental attention heads (see Appendix G); and (2) the inherent limitations of CLIP models. Tasks such as CR (code reasoning), NC (numerical calculation), and TT (text translation) demand beyond visual perception, but understanding and reasoning capabilities that CLIP is not explicitly trained for.
@@
Table 6: Comparison on COCO-CN for retrieval.
@@
Table 7: Comparison on Flickr30k-CNA for retrieval.
@@
Table 8: Top-1 accuracy on the ImageNet-1k val set for zero-shot classification, and mean-R on the test sets of MS COCO and Flickr30k for text-to-image retrieval.
5.2.2 AAT-improved Chinese-CLIP for Retrieval
To further validate the versatility of AAT, we evaluate it on non-English linguistic CLIP models. Following the same evaluation protocol as with CLIP, we assess AAT-improved Chinese-CLIP on COCO-CN and Flickr30k-CNA, comparing against their vanilla counterparts.
COCO-CN
The text-to-image retrieval results are shown in Table 6. On the test set, AAT achieves a 1.9$\sim$3.2higher mean-R across model sizes using both GA and BP. On the all set, the gain is even more pronounced, reaching up to 6.4%. For image-to-text retrieval, AAT delivers a 2.5$\sim$6.7improvement on the test set and 3.9$\sim$11.1on the all set. These advances surpass those observed in CLIP, suggesting its effectiveness is correlated with the quality of the original representations.
Flickr30k-CNA
As shown in Table 7, AAT improves mean-R by 2.21$\sim$3.2on the test set and 3.14$\sim$4.33on the all set for text-to-image retrieval. For image-to-text retrieval, it outperforms the baseline by around 3% on the test set and up to 8.98% on the all set. Consistent with the results on COCO-CN, AAT delivers greater improvements for Chinese-CLIP compared to CLIP.
5.2.3 AAT-improved CLIP for Zero-shot Classification
@@
We evaluate the resulting models on the ImageNet-1k val set. As shown in Table 8, they consistently achieve 1$\sim$1.5gains in top-1 accuracy. To further verify AAT’s generalization, we test retrieval performance of these models on MS COCO and Flickr30k. The models continue to deliver strong improvements, with no noticeable degradation compared to counterparts optimized with the original $\mathcal{D}$ (see Table 3 and 4), with differences of less than 0.6% on MS COCO and 0.3% on Flickr30k.
Figure 3: Mean-R on the COCO-CN all set vs. $\beta$ values in AAT-GA, using the ViT-B-based model.
@@
Table 9: Mean-R vs. the size of $\mathcal{D}$: AAT-improved CLIP is evaluated on the MS COCO all set, and AAT-improved Chinese-CLIP on the COCO-CN all set, both using ViT-B-based models.
@@
Table 10: Recall rates on Flickr30k-CNA for text-to-image retrieval, based on ViT-B. AAT is optimized on the COCO-CN val set and a 1k-sampled subset of the Flickr30k-CNA val set, respectively.
5.3 Ablation Study
5.3.1 The Value of $\beta$
We empirically set $\beta=0.1$ in AAT-GA and study its impact on performance. Figure 3 shows the relationship between decreasing $\beta$ values and mean-R on the COCO-CN all set. As $\beta$ decreases from 1 to 0.02, mean-R initially improves, peaking at $\beta=0.1$, and then declines. For AAT-BP, we examine the $\beta$ values derived from the sigmoid function applied to the learnable $\alpha$ parameters with temperature $\tau$. We observe that $\beta$ values in deeper transformer layers tend to be more polarized—clustered near 0 or 1—whereas those in shallower layers remain closer to the initial value of 1. This suggests that gradient propagation during BP optimization varies across layers, which may help explain the performance differences between AAT-GA and AAT-BP observed in certain scenarios.
5.3.2 The Size of $\mathcal{D}$
We ablate the effect of the validation set $\mathcal{D}$ size, which originally contains 1k image-text pairs for AAT optimization. Subsets of 100, 200, and 500 image-text pairs are randomly sampled from $\mathcal{D}$, and the resulting mean-R for AAT-improved CLIP and Chinese-CLIP are reported in Table 9. Even with only 100 image-text pairs, AAT outperforms the baseline ($\text{size}=0$) by around 1% for CLIP and 3% for Chinese-CLIP. As the size increases to 500, the further performance gain becomes marginal.
5.3.3 Selection of AAT Optimization Data
As described in Section 5.1, we use a 1k-subset from MS COCO or COCO-CN to optimize AAT. To examine the impact of optimization data selection and potential domain bias, we construct a new $\mathcal{D}$ using 1k images from the Flickr30k-CNA val set, each paired with one corresponding caption. We then re-run AAT using this in-domain $\mathcal{D}$ and compare its performance on Flickr30k-CNA against that of AAT optimized with COCO-CN samples. As shown in Table 10, using in-domain data yields only marginal improvements (typically $<$0.5%), implicitly demonstrating the strong generalization of AAT-refined representations across domains.
size metric vanilla AAT-BP AAT-GA AAT-GA (optimized)
ViT-B mean-R 58.9 60.9 61.5 61.2
runtime - 2 min (0.3 GPU hours) 30 min (4 GPU hours) 18 min (2.4 GPU hours)
ViT-L mean-R 63.7 66.2 66.1 65.9
runtime - 5 min (0.7 GPU hours) 1.5 h (12 GPU hours) 27 min (3.6 GPU hours)
ViT-H mean-R 66.8 68.1 68.7 68.3
runtime - 8 min (1.1 GPU hours) 3 h (24 GPU hours) 50 min (6.7 GPU hours)
@@ 5.3.3 Selection of AAT Optimization Data
Table 11: Mean-R on the MS COCO test set for text-to-image retrieval vs. the actual AAT optimization runtime (GPU hours). “Optimized” refers to AAT-GA with optimized hyper-parameter settings.
5.4 AAT Optimization Runtime
We use 8×8$\times$8 × NVIDIA RTX 4090 GPUs and an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz to optimize AAT. To demonstrate its real-world feasibility, we report the actual runtime and GPU hours under the standard configuration described in Section 5.1, as listed in the fourth and fifth columns of Table 11. While GA offers key advantages for low-precision (like INT8/INT16) inference devices, as no need for gradients during optimization, it incurs roughly 20×20$\times$20 × the time cost of BP due to the large number of fitness evaluations (even with the small validation set $\mathcal{D}$).
AAT is intentionally designed as an “out-of-the-box” practical solution, with GA hyper-parameters, such as the number of generations and crossover/mutation rates, kept consistent across different model sizes and languages to minimize manual tuning. Despite its simplicity and demonstrated effectiveness, the use of unified settings may introduce redundant computational overhead. Further hyper-parameter tuning could significantly reduce this cost without compromising performance.
Specifically, we implement the following optimizations. First, based on the ablated head analysis in Appendix C, where effective ablation ratios typically fall between 30% and 40%, we skip fitness evaluations for candidates with ablation ratios exceeding 60%. Second, we reduce the GA population size for ViT-L-based and ViT-H-based models, as smaller populations maintain comparable performance while significantly lowering runtime. Lastly, running 100 generations proves excessive: performance gains plateau during the final 30% of generations, even as fitness scores continue to improve—suggesting that over-optimization on $\mathcal{D}$ offers diminishing benefit on test data.
With these targeted adjustments, we achieve a significantly more efficient GA setup, reducing optimization overhead by 40%, 70%, and 72% for the base, large, and huge models, respectively, with only a slight drop in performance, as shown in Table 11. Detailed performance and runtime comparisons under different hyper-parameter configurations are provided in Appendix D. Notably, as a post-training technique, both AAT-BP and AAT-GA incur minimal overhead compared to training CLIP from scratch, underscoring their practicality for real-world deployment.

### v2

@@ 4.3 AAT with Back-Propagation
Table 4: Comparison on Flickr30k for cross-modal retrieval. The symbols in the table have the same meaning as those in Table 3.
@@
tasks ACT CEL COL CNT FIG OBJ OCR POS SCE mean CR NC $T$ mean
@@
Table 5: Text-to-image retrieval results on ReCoS using ViT-B, evaluated with R@1 as the metric.
5 Experiments
5.1 Experimental Settings
5.1.1 Models
We evaluate AAT using CLIP [radford2021learning] and Chinese-CLIP [yang2022chinese], across ViT-B, ViT-L, and ViT-H. Specifically, we use OpenCLIP released models [ilharco] trained on LAION-2B (a subset of LAION-5B [schuhmann2022laion]) and Chinese-CLIP released models trained on 200M Chinese image-text pairs.
5.1.2 AAT Optimization Data
5.1.3 AAT-GA Implementation
The key hyper-parameters for GA are configured as follows: the population size is determined by the number of heads—48 for ViT-B, 96 for ViT-L, and 128 for ViT-H. A two-point crossover is applied with a probability of 0.9, while mutation occurs with a probability of 0.5 using the flip-bit strategy. The tournament selection with a size of 3 is employed. The evolution proceeds for up to 100 generations, with early stopping triggered when fitness gains plateau or population diversity becomes low.
5.1.4 AAT-BP Implementation
The BP optimization settings are as follows: learning rate of 2e-2 for CLIP and 5e-2 for Chinese-CLIP, with a batch size of 256. ViT-B models are trained for 32 epochs, and larger models for 64. All other hyper-parameters follow the OpenCLIP codebase [ilharco].
5.1.5 Benchmarks
We evaluate AAT-improved CLIP upon MS COCO [lin2014microsoft], Flickr30k [young2014image], and ReCoS [chen2024recos], and assess AAT-improved Chinese-CLIP on two widely-used Chinese retrieval benchmarks: COCO-CN [li2019coco] and Flickr30k-CNA [xie2023ccmb]. For MS COCO, Flickr30k, COCO-CN, Flickr30k-CNA, we report results on two splits: the test split (the original test set) and the all split (all samples in the dataset), enabling a comprehensive large-scale retrieval analysis. We further evaluate zero-shot classification on the ImageNet-1k val set [deng2009imagenet]. Dataset and metric details are provided in Appendix 7.
5.1.6 Baseline
Aside from vanilla CLIP, we also compare AAT to two recent strong PEFT baselines, namely CLIP-LoRA [zanella2024low] and ReFT [wu2024reft]. Detailed settings for both are as follows.
Vanilla CLIP
The vanilla CLIP and Chinese-CLIP models with their original, unmodified weights (prior to any improvement via AAT) serve as our primary baselines.
CLIP-LoRA
For fairness, we apply LoRA (using rank 1 or 2) to the Q/K projections of each head from image encoder, mirroring AAT’s intervention points. To better match AAT-BP’s parameter count, we also include a light-weight variant (LoRA-lite) that targets only the final image encoder layer. We adopt CLIP-LoRA’s training setup: learning rate 2e-4, batch size 32, and 64 epochs (chosen via search).
ReFT
Following the paper, we use rank 4 for the low-rank projection $R$. A preparatory search over intervention positions shows that applying ReFT to all tokens slightly outperforms restricting it to the class token or the top-5 tokens for CLIP (contrary to some LLM findings). For completeness, we implement both DiReFT and LoReFT across all image-encoder layers, and parameter-reduced variants (DiReFT- / LoReFT-lite) applied only to the last layer. We train with learning rate 2e-5 and batch size 32 (higher rates tend to collapse $R$), for 96 epochs until accuracy saturates.
5.2 Compared with Vanilla CLIP
5.2.1 Cross-modal Retrieval
@@
[…]
@@
[…]
@@
Table 6: Comparison on COCO-CN for cross-modal retrieval. The symbols in the table have the same meaning as those in Table 3.
@@
Table 7: Comparison on Flickr30k-CNA for cross-modal retrieval. The symbols in the table have the same meaning as those in Table 3.
ReCoS
To evaluate broader effectiveness, Table 5 reports text-to-image retrieval on the challenging ReCoS, spanning 12 subsets across diverse domains (subset abbreviations in Appendix 7). We group subsets by image domain: natural (e.g., landscapes, animals, movie scenes, profile photos) and non-natural (e.g., code snippets, characters, formulas) that are typically synthetic with uniform backgrounds.
AAT consistently outperforms on natural subsets, with a mean R@1 gain of around 3%. However, its performance declines on non-natural images. We attribute this drop to two main factors: (1) significant domain shifts between the test images and those used to identify detrimental attention heads (see Appendix 13); and (2) the inherent limitations of CLIP models. Tasks such as CR (code reasoning), NC (numerical calculation), and TT (text translation) demand beyond visual perception, but understanding and reasoning capabilities that CLIP is not explicitly trained for.
COCO-CN
[…]
Flickr30k-CNA
[…]
ViT-B ViT-L ViT-H
#param MS COCO Flickr30k Avg #param MS COCO Flickr30k Avg #param MS COCO Flickr30k Avg
T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T
vanilla n/a 58.9 76.0 82.98 94.30 78.05 n/a 63.7 77.0 86.91 94.90 80.63 n/a 66.8 78.7 88.65 95.63 82.45
AAT-GA 0 61.5 77.4 84.91 94.40 79.55 0 66.1 79.1 88.13 95.33 82.17 0 68.7 80.4 89.76 96.30 83.79
AAT-BP 144 60.9 77.0 84.60 94.33 79.21 384 66.2 77.8 88.43 95.00 81.86 512 68.1 80.4 89.73 96.40 83.66
LoRA-lite 3.1k 59.2 76.3 83.23 94.10 78.21 4.1k 64.2 77.4 87.28 94.60 80.87 5.1k 67.2 79.0 88.83 95.60 82.66
LoRA (r=1) 36.9k 62.2 77.2 85.39 93.63 79.61 98.3k 65.8 78.7 88.04 95.27 81.95 163.8k 68.7 79.9 89.93 96.10 83.66
LoRA (r=2) 73.8k 62.7 77.5 85.73 93.98 79.98 196.6k 66.4 78.6 88.29 95.22 82.13 327.5k 69.1 80.3 89.85 96.40 83.91
DiReFT-lite 6.1k 60.1 75.7 83.82 93.83 78.36 8.2k 64.3 76.9 87.32 94.43 80.74 10.2k 67.1 78.9 88.91 95.90 82.70
DiReFT 73.8k 62.0 76.3 85.18 93.50 79.25 196.6k 65.7 77.4 88.16 94.50 81.44 327.5k 68.9 78.6 89.95 95.50 83.24
LoReFT-lite 6.1k 59.9 75.2 83.49 93.73 78.08 8.2k 64.3 77.1 87.11 94.53 80.76 10.2k 67.0 78.4 88.67 94.27 82.09
LoReFT 73.8k 61.1 75.2 84.70 93.93 78.73 196.6k 65.4 76.7 87.76 94.10 80.99 327.5k 67.8 78.3 89.47 95.33 82.73
Table 8: Mean-R comparison for text-to-image (T2I) and image-to-text (I2T) retrieval across base, large, and huge model variants.
@@
Table 9: Top-1 accuracy on ImageNet-1k val set for zero-shot classification, and mean-R on MS COCO and Flickr30k test sets.
5.2.2 Zero-shot Classification
@@
We test the resulting models on the ImageNet-1k val set and observe consistent top-1 accuracy gains of $\sim1$.5% (Table 9). To further assess generalization, we test retrieval on MS COCO and Flickr30k. Performance remains strong, with no noticeable degradation relative to models optimized on the original $\mathcal{D}$ (Tables 3 and 4); text-to-image differences are under 0.6% on MS COCO and 0.3% on Flickr30k.
5.3 Compared with PEFT
Table 8 reports recall for text-to-image and image-to-text retrieval on MS COCO and Flickr30k, along with the number of additional learnable parameters.
5.3.1 AAT vs. LoRA.
[…]
5.3.2 AAT vs. ReFT.
[…]
5.4 Ablation Study
5.4.1 The Value of $\beta$
We empirically set $\beta=0.1$ in AAT-GA and examine its effect on performance. Figure 3 plots mean-R on the COCO-CN all set as $\beta$ decreases from 1 to 0.02: performance rises, peaks at $\beta=0.1$, then declines. For AAT-BP, we analyze the head-specific $\beta$ values obtained by applying a sigmoid (with temperature $\tau$) to the learnable $\alpha$ parameters. Deeper layers exhibit more polarized $\beta$ (near 0 or 1), whereas shallower layers remain closer to the initial value of 1. This suggests depth-dependent gradient propagation during BP, which may account for the result differences between AAT-GA and AAT-BP in certain cases.
Figure 3: Mean-R for text-to-image retrieval on the COCO-CN all set vs. $\beta$ values in AAT-GA, using the ViT-B-based model.
5.4.2 The Size of $D$
We study the effect of the validation set size for AAT optimization. Starting from $\mathcal{D}$ with 1k image–text pairs, we randomly sample subsets of 500, 200, and 100 pairs and report mean-R for both CLIP and Chinese-CLIP in Table 10. Even with only 100 pairs, AAT surpasses the baseline ($\text{size}=0$) by 1% on CLIP and 3% on Chinese-CLIP. As increasing the size to 500, it yields only marginal additional gains.
5.4.3 Selection of AAT Optimization Data
As described in Section 5.1.2, we optimize AAT on a 1k subset from MS COCO or COCO-CN. To examine the effect of used data and potential domain bias, we construct an in-domain $\mathcal{D}$ using 1k Flickr30k-CNA val image-text pairs. We then re-run AAT with this in-domain set and compare Flickr30k-CNA performance to AAT optimized on COCO-CN. As shown in Table 11, using in-domain data yields only marginal gains (typically $0.5\%$), indicating strong cross-domain generalization of AAT-refined representations.
5.5 AAT Optimization Runtime
We optimize AAT on $8{\times}$ NVIDIA RTX 4090 GPUs with an Intel(R) Xeon(R) Platinum 8358P @ 2.60GHz platform. To demonstrate real-world feasibility, we list wall-clock runtime and GPU-hours under the standard setup in Section 5.1 (see the 4th–5th columns of Table 13). While GA is attractive for low-precision (INT8/INT16) inference hardware, as requiring no gradients, it is roughly $20{\times}$ slower than BP due to a vast number of fitness evaluations, even with a small $\mathcal{D}$.
AAT is designed as an “out-of-the-box” solution where GA hyper-parameters (e.g., generations, crossover/mutation rates) are kept fixed across model sizes to minimize manual adjustment. This simplicity introduces redundant overhead. We therefore apply the following practical optimizations:
  1. 1.
    Skipping: Based on the analysis in Appendix 9, where effective ablation ratios typically fall between 30% and 40%, we skip candidates whose ratios exceed 60%.
  2. 2.
    Reducing population size for ViT-L/H: We find that the reduced population sizes maintain comparable accuracy while significantly cutting runtime.
  3. 3.
    Fewer generations: We find that performance plateaus in the final 30% of generations, despite a continued fitness gain, suggesting that over-optimization on $\mathcal{D}$ offers diminishing benefit on test data.
These adjustments yield substantial overhead savings—reducing GA runtime by 40%, 70%, and 72% for the base, large, and huge models, with only minor performance drops (Table 13; detailed comparisons in Appendix 10). As post-training technique, both AAT-BP and AAT-GA impose minimal cost compared to training CLIP from scratch, underscoring their practicality for real-world deployment.
@@ 5.5 AAT Optimization Runtime
Table 10: Mean-R vs. the size of $\mathcal{D}$: CLIP is evaluated on the MS COCO all set, and Chinese-CLIP on the COCO-CN all set, both using ViT-B-based models, for text-to-image retrieval.
@@
Table 11: Flickr30k-CNA text-to-image retrieval results based on ViT-B. CO indicates that AAT is optimized on the COCO-CN val set, while F30k indicates optimization on a 1k-sample subset of the Flickr30k-CNA val set.
tasks ACT CEL COL CNT FIG OBJ OCR POS SCE mean
$\mathcal{S}_{\text{gain}}$ 2.26 2.91 2.02 2.59 2.29 2.84 3.79 2.34 2.40 2.60
$\mathcal{S}_{\text{loss}}$ 2.60 3.87 2.74 2.77 2.72 3.05 4.03 2.74 3.25 3.09
Table 12: Average text-related importance of ablated heads in the last 4 layers on natural ReCoS (reported values in units of 1e-2).
5.6 An Interpretability Perspective on AAT
We investigate the interpretability of the ablated heads using TextSpan [gandelsmaninterpreting] to correlate head-level effects to linguistic semantics. Using a ViT-B-based CLIP model, for each image–text pair we extract the text and image embeddings and, following TextSpan, decompose the image embedding into 144 ($\times12$) head-specific image embeddings. Each one is then multiplied by the text embedding to yield a $\times12$ matrix of text-related importance, where larger values indicate higher semantic relevance. This matrix reveals which heads in vanilla CLIP are most (or least) aligned with the text.
We probe these semantics on the natural subsets of ReCoS. Specifically, we partition all natural pairs into those that flip from incorrect to correct after applying AAT-GA ($\mathcal{S}_{\text{gain}}$) and those that flip from correct to incorrect ($\mathcal{S}_{\text{loss}}$). For each pair, we compute the cumulative text-related importance (from the $\times12$ matrix) of the heads ablated by AAT in the last four layers, as recommended by TextSpan. As shown in Table 12, the ablated heads in $\mathcal{S}_{\text{gain}}$ have lower average importance, whereas those in $\mathcal{S}_{\text{loss}}$ have higher importance. This supports the view that suppressing less text-relevant heads tends to improve image–text alignment (and suppressing highly text-relevant heads can harm it) [gandelsmaninterpreting].
These findings complement the CLIP interpretability research [gandelsmaninterpreting], which manually removes heads associated with spurious cues in specific layers and observes improved representations. Instead, AAT offers an automated, data-driven scheme that can target detrimental heads across all layers. Encouragingly, AAT naturally focuses on heads with lower text-related importance, echoing those manual intervention results. It thereby reinforces the conclusion that attention heads exhibit meaningful semantic structure.
size metric vanilla BP GA GA (opt)
B mR 58.9 60.9 61.5 61.2
runtime - 2 min (0.3 GHs) 30 min (4 GHs) 18 min (2.4 GHs)
L mR 63.7 66.2 66.1 65.9
runtime - 5 min (0.7 GHs) 1.5 h (12 GHs) 27 min (3.6 GHs)
H mR 66.8 68.1 68.7 68.3
runtime - 8 min (1.1 GHs) 3 h (24 GHs) 50 min (6.7 GHs)
@@ 5.6 An Interpretability Perspective on AAT
Table 13: Mean-R on MS COCO test set for text-to-image retrieval vs. actual optimization runtime. “Opt” denotes AAT-GA with optimized hyper-parameter settings; “GHs” denotes the GPU hours.